# Conditional Polarization Guidance for Camouflaged Object Detection

### 31 Mar 2026 v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026
@@ 1. Introduction
Recent RGB-polarization COD methods have shown encouraging results, confirming the value of polarization in reducing RGB ambiguity. For example , IPNet(9) designs cross-modal modules to effectively integrate polarization and RGB cues at multiple levels, improving contrast in complex scenes. Similarly, HIPFNet (Wang et al., 2025) specifically designed a polarization image enhancer that accepts Stokes parameters to optimize the unique information contained in orthogonal polarization images. Although these methods have achieved promising performance, most existing RGB–polarization approaches rely on symmetric dual-stream encoding or heavy multimodal fusion, which increases model complexity while underexploring how polarization can explicitly guide hierarchical RGB representation learning.
To this end, this paper proposes a novel RGB–polarization architecture for COD, dubbed CPGNet, which leverages polarization as conditional structural guidance for RGB feature learning. Specifically, CPGNet is composed of three coordinated components. First, a Polarization Integration Module constructs reliable polarization guidance by exploiting the complementarity of DoLP and AoLP while suppressing less informative responses. Building upon this, a Polarization Guidance Flow progressively injects the derived polarization guidance into hierarchical RGB features, enabling the network to focus on subtle appearance discrepancies; meanwhile, an edge-guided frequency refinement strategy further enhances discriminative high-frequency cues to break camouflage patterns. Finally, an Iterative Feedback Decoder IFD performs coarse-to-fine prediction refinement through hierarchical feedback calibration, improving localization accuracy and boundary completeness in ambiguous camouflage regions. Together, these components establish a unified conditional guidance framework in which polarization continuously regulates RGB representation learning rather than being directly fused as a parallel modality. Experiments on PCOD-1200 show that CPGNet performs favorably against recent RGB-only and multimodal methods. In addition, results on RGBP-Glass suggest that the proposed framework may also generalize to other RGB-polarization tasks.
@@ 2.1. Polarization-based Vision
(2) $DoLP=\frac{\sqrt{S_{1}^{2}+S_{2}^{2}}}{S_{0}},AoLP=\frac{1}{2}\mathrm{arc}\tan\left(\frac{S_{2}}{S_{1}}\right),$
@@ 3.3.1. Polarization Guidance Enhance
3.3.2. Edge guided Frequency Module
Previous studies have proved that the amplitude and phase component obtained by the Fourier transform contains more critical information for the object. In particular,the polarization edge structures tend to exhibit stronger responses in the amplitude spectrum. Based on this, we leverage polarization-derived edge priors to guide frequency-domain feature enhancement. We apply the Fast Fourier Transform (FFT) to the RGB and edge features and perform operations on their amplitude components. After this operation, we transform again to the space domain with the Inverse Fast Fourier Transform. This allows polarization-derived edges to provide reliable structural guidance during frequency enhancement. This can be specified more precisely by the following equations:
@@ 3.4. Iterative Feedback Decoder
Humans often identify camouflaged objects through repeated observation, and the perception process typically begins with a coarse understanding of the scene and gradually focuses on regions of interest to refine the recognition of subtle details. This process reflects the collaborative interaction between bottom-up and top-down perceptual mechanisms. Inspired by this, we propose an Iterative Feedback Decoder , as illustrated in Figure 2 (a). The iterative process progressively refines predictions by leveraging feedback from previous stages, enabling coarse-to-fine optimization. This process can be formulated by the following equations:
@@ 3.5. Loss Function
To effectively guide each iterative stage, we apply supervision to the prediction results after each iteration of the decoder. Specifically, we generate intermediate predictions $P^{(i)}$ from each iteration (with the total number of iterations set to 2) and $P_{4}$ denotes the coarse prediction generated from the last stage. Following (Fan et al., 2020a), we employ weighted binary cross-entropy loss $\mathcal{L}_{\mathrm{\omega BCE}}$ and weighted intersection-over-union loss $\mathcal{L}_{\mathrm{\omega IoU}}$ to better supervise hard-to-detect pixels, combined as $\mathcal{L}_{str}=\mathcal{L}_{\mathrm{\omega IoU}}+\mathcal{L}_{\mathrm{\omega BCE}}$. This explicit supervision ensures that both high-level semantic localization and low-level boundary refinement are jointly optimized. Overall, our total loss function is formulated as:
@@ 4.1. Experimental Setup
ZoomNetxt (Pang et al., 2024a) TPAMI’24 0.945 0.974 0.006 0.909
@@
4.2. Qualitative and Quantitative Evaluation
Quantitative Analysis. We quantitatively compare our method with 20 state-of-the-art (SOTA) methods on the PCOD dataset, including 12 RGB-based methods and 8 multimodal segmentation methods. As shown in Table 1, our method achieves the best overall performance. As shown in Table 1, our method achieves the best performance across all evaluation metrics, with scores of 0.951 in $S_{\alpha}$, 0.923 in $F_{\beta}^{w}$, 0.981 in $E_{\phi}$, and 0.005 in $M$. Our method consistently outperforms all SAM-based methods, demonstrating its superiority in capturing challenging object characteristics. Compared with the second-best model HIPFNet (Wang et al., 2025), our method achieves a significant improvement of 0.007 in $S_{\alpha}$. The strong $S_{\alpha}$ result further indicates that the proposed conditional polarization guidance effectively preserves informative structural cues for camouflaged object detection.
Qualitative Analysis. Figure 3 presents a visual comparison between our proposed method and other advanced methods. We evaluate challenging camouflaged images featuring different object types under extreme scenes involving multiple targets and small objects. Our approach produces cleaner and more complete predictions, especially in cluttered backgrounds with strong texture interference. These results indicate that our model effectively captures rich object details. AoLP and DoLP provide effective auxiliary cues due to their sensitivity to material properties, thereby improving the discriminability of camouflaged targets. A closer inspection of the visual results further reveals that the advantage of our method becomes more evident in scenes with complex textures, weak object boundaries, and severe appearance ambiguity. In these cases, several competing methods either miss object parts or produce scattered false positives in the background. By contrast, our method achieves more reliable localization in these challenging cases. This observation suggests that the proposed conditional guidance mechanism helps the network better exploit the complementary structural and material cues provided by polarization.
Comparisons on transparent object detection datasets. To further assess the generality of the proposed framework, we additionally evaluate it on the RGBP-Glass dataset, a related RGB-polarization glass segmentation benchmark. We quantitatively compare our method with 8 methods on the RGBP-Glass dataset. As shown in Table 2, our method achieves the best performance on $M$, IoUIoU, and $BER$, with scores of 0.082, 85.09, and 7.84, respectively, while also obtaining competitive results on other metrics. This suggests that the proposed conditional polarization guidance generalizes well to other RGB-polarization dense prediction tasks. As shown in Fig. 4, our method achieves accurate segmentation of transparent objects under challenging conditions, such as complex backgrounds or varying illumination.
Comparisons on Non-polarization Datasets To further assess the generalization capability of our framework, we evaluate it on widely used COD benchmarks, including CAMO, COD10K, and NC4K. As these datasets do not provide polarization measurements, we approximate polarization cues using surrogate signals, where Sobel-based RGB gradients and estimated depth maps are adopted to mimic DoLP and AoLP, respectively. As shown in Table 3, our method performs favorably against other competitive COD methods on these datasets. As shown in Fig. 5, our method demonstrates superior performance in challenging scenarios, particularly exhibiting more accurate boundary prediction for camouflaged objects. While our method is primarily designed for polarization-driven perception and is not specifically optimized for RGB-only settings, it still achieves competitive performance with a relatively low number of parameters, demonstrating the robustness and transferability of the proposed polarization-guided design.
@@ 4.2. Qualitative and Quantitative Evaluation
4.3. Ablation Study
Impact of Polarization Cues.
As shown in Table 4, incorporating polarization cues consistently improves the performance over the RGB-only baseline. Compared with baseline (A), introducing AoLP (B) leads to gains in structure-related metrics, including $S_{\alpha}$ and $E_{\phi}$, indicating that AoLP provides useful directional and structural information for camouflage perception. Introducing DoLP alone (C) further improves the results, suggesting that DoLP also provides effective complementary cues for distinguishing camouflaged targets. These observations indicate that both AoLP and DoLP are beneficial, while capturing different aspects of polarization information. However, each cue alone can only provide partial polarization evidence. This motivates our design to jointly model them for constructing more reliable polarization guidance, and supports our design philosophy of treating polarization as structural guidance rather than direct feature fusion.
@@ 4.3. Ablation Study
Effectiveness of the Iterative Feedback Decoder. Finally, removing the Iterative Feedback Decoder (IFD) (I) results in the most noticeable performance drop, especially in $F_{\beta}^{w}$ and $S_{\alpha}$, indicating that coarse-to-fine feedback refinement is crucial for recovering ambiguous structures and improving boundary completeness.
Overall Discussion.
Overall, all ablation variants perform worse than the full model, which confirms that the superiority of CPGNet comes from the coordinated effect of PIM, PGF, and IFD, where PGF further benefits from the joint contribution of PGE and EFM. This also suggests that effective camouflage perception requires not only reliable polarization guidance, but also its progressive propagation and refinement throughout the network.
@@ 5. Limitation
The proposed CPGNet still has several limitations. Its performance depends on the quality of polarization measurements, and unstable DoLP/AoLP responses in weak-polarization regions may affect the reliability of the derived guidance cues. Figure 7 shows that RGB lacks discriminative features, while both polarization and depth modalities offer limited complementary cues, making it challenging for many methods to detect camouflaged targets.

### 01 Apr 2026 v2

@@ Abstract.
camouflaged object detection, polarization, RGB-polarization, multimodal learning, image segmentation
copyright: nonejournalyear: 2026
@@ 1. Introduction
Recent RGB-polarization COD methods have shown encouraging results, confirming the value of polarization in reducing RGB ambiguity. For example, IPNet (9) designs cross-modal modules to effectively integrate polarization and RGB cues at multiple levels, improving contrast in complex scenes. Similarly, HIPFNet (Wang et al., 2025) specifically designed a polarization image enhancer that accepts Stokes parameters to optimize the unique information contained in orthogonal polarization images. Although these methods have achieved promising performance, most existing RGB–polarization approaches, as illustrated in Fig. 1, rely on symmetric dual-stream encoding or heavy multimodal fusion, which increases model complexity while underexploring how polarization can explicitly guide hierarchical RGB representation learning.
To this end, this paper proposes a novel RGB–polarization architecture for COD, dubbed CPGNet, which leverages polarization as conditional structural guidance for RGB feature learning. Specifically, CPGNet is composed of three coordinated components. First, a Polarization Integration Module constructs reliable polarization guidance by exploiting the complementarity of DoLP and AoLP while suppressing less informative responses. Building upon this, a Polarization Guidance Flow progressively injects the derived polarization guidance into hierarchical RGB features, enabling the network to focus on subtle appearance discrepancies; meanwhile, an edge-guided frequency refinement strategy further enhances discriminative high-frequency cues to break camouflage patterns. Finally, an Iterative Feedback Decoder (IFD) performs coarse-to-fine prediction refinement through hierarchical feedback calibration, improving localization accuracy and boundary completeness in ambiguous camouflage regions. Together, these components establish a unified conditional guidance framework in which polarization continuously regulates RGB representation learning rather than being directly fused as a parallel modality. Experiments on PCOD-1200 show that CPGNet performs favorably against recent RGB-only and multimodal methods. In addition, results on RGBP-Glass suggest that the proposed framework may also generalize to other RGB-polarization tasks.
@@ 2.1. Polarization-based Vision
(2) $\mathrm{DoLP}=\frac{\sqrt{S_{1}^{2}+S_{2}^{2}}}{S_{0}},\quad\mathrm{AoLP}=\frac{1}{2}\arctan\left(\frac{S_{2}}{S_{1}}\right),$
@@ 3.3.1. Polarization Guidance Enhance
3.3.2. Edge-Guided Frequency Module
Previous studies have proved that the amplitude and phase components obtained by the Fourier transform contain more critical information for the object. In particular, the polarization edge structures tend to exhibit stronger responses in the amplitude spectrum. Based on this, we leverage polarization-derived edge priors to guide frequency-domain feature enhancement. We apply the Fast Fourier Transform (FFT) to the RGB and edge features and perform operations on their amplitude components. After this operation, we transform again to the space domain with the Inverse Fast Fourier Transform. This allows polarization-derived edges to provide reliable structural guidance during frequency enhancement. This can be specified more precisely by the following equations:
@@ 3.4. Iterative Feedback Decoder
Humans often identify camouflaged objects through repeated observation, and the perception process typically begins with a coarse understanding of the scene and gradually focuses on regions of interest to refine the recognition of subtle details. This process reflects the collaborative interaction between bottom-up and top-down perceptual mechanisms. Inspired by this, we propose an Iterative Feedback Decoder, as illustrated in Fig. 2(a). The iterative process progressively refines predictions by leveraging feedback from previous stages, enabling coarse-to-fine optimization. This process can be formulated by the following equations:
@@ 3.5. Loss Function
To effectively guide each iterative stage, we apply supervision to the prediction results after each iteration of the decoder. Specifically, we generate intermediate predictions $P^{(i)}$ from each iteration (with the total number of iterations set to 2) and $P_{4}$ denotes the coarse prediction generated from the last stage. Following (Fan et al., 2020a), we employ weighted binary cross-entropy loss $\mathcal{L}_{\mathrm{\omega BCE}}$ and weighted intersection-over-union loss $\mathcal{L}_{\mathrm{\omega IoU}}$ to better supervise hard-to-detect pixels, combined as $\mathcal{L}_{\mathrm{str}}=\mathcal{L}_{\mathrm{\omega IoU}}+\mathcal{L}_{\mathrm{\omega BCE}}$. This explicit supervision ensures that both high-level semantic localization and low-level boundary refinement are jointly optimized. Overall, our total loss function is formulated as:
@@ 4.1. Experimental Setup
ZoomNeXt (Pang et al., 2024a) TPAMI’24 0.945 0.974 0.006 0.909
@@
4.2. Qualitative and Quantitative Evaluation
Quantitative Analysis. We quantitatively compare our method with 20 state-of-the-art (SOTA) methods on the PCOD-1200 dataset, including 12 RGB-based methods and 8 multimodal segmentation methods. As shown in Tab. 1, our method achieves the best performance across all evaluation metrics, with scores of 0.951 in $S_{\alpha}$, 0.923 in $F_{\beta}^{w}$, 0.981 in $E_{\phi}$, and 0.005 in $M$. Our method consistently outperforms all SAM-based methods, demonstrating its superiority in capturing challenging object characteristics. Compared with the second-best model HIPFNet (Wang et al., 2025), our method achieves a significant improvement of 0.007 in $S_{\alpha}$. The strong $S_{\alpha}$ result further indicates that the proposed conditional polarization guidance effectively preserves informative structural cues for camouflaged object detection.
@@ 4.2. Qualitative and Quantitative Evaluation
Qualitative Analysis. Fig. 3 presents a visual comparison between our proposed method and other advanced methods. We evaluate challenging camouflaged images featuring different object types under extreme scenes involving multiple targets and small objects. Our approach produces cleaner and more complete predictions, especially in cluttered backgrounds with strong texture interference. These results indicate that our model effectively captures rich object details. AoLP and DoLP provide effective auxiliary cues due to their sensitivity to material properties, thereby improving the discriminability of camouflaged targets. A closer inspection of the visual results further reveals that the advantage of our method becomes more evident in scenes with complex textures, weak object boundaries, and severe appearance ambiguity. In these cases, several competing methods either miss object parts or produce scattered false positives in the background. By contrast, our method achieves more reliable localization in these challenging cases. This observation suggests that the proposed conditional guidance mechanism helps the network better exploit the complementary structural and material cues provided by polarization.
@@
Comparisons on Transparent Object Detection Datasets. To further assess the generality of the proposed framework, we additionally evaluate it on the RGBP-Glass dataset, a related RGB-polarization glass segmentation benchmark. We quantitatively compare our method with 8 methods on the RGBP-glass dataset. As shown in Tab. 2, our method achieves the best performance on $M$, IoUIoU, and $BER$, with scores of 0.082, 85.09, and 7.84, respectively, while also obtaining competitive results on other metrics. This suggests that the proposed conditional polarization guidance can generalize well to other RGB-polarization dense prediction tasks. As shown in Fig. 4, our method achieves accurate segmentation of transparent objects under challenging conditions, such as complex backgrounds or varying illumination.
Comparisons on Non-polarization Datasets. To further assess the generalization capability of our framework, we evaluate it on widely used COD benchmarks, including CAMO, COD10K, and NC4K. As these datasets do not provide polarization measurements, we approximate polarization cues using surrogate signals, where Sobel-based RGB gradients and estimated depth maps are adopted to mimic DoLP and AoLP, respectively. As shown in Tab. 3, our method performs favorably against other competitive COD methods on these datasets. As shown in Fig. 5, our method demonstrates superior performance in challenging scenarios, particularly exhibiting more accurate boundary prediction for camouflaged objects. While our method is primarily designed for polarization-driven perception and is not specifically optimized for RGB-only settings, it still achieves competitive performance with a relatively low number of parameters, demonstrating the robustness and transferability of the proposed polarization-guided design.
4.3. Ablation Study
@@
Impact of Polarization Cues. As shown in Tab. 4, incorporating polarization cues consistently improves the performance over the RGB-only baseline. Compared with baseline (A), introducing AoLP (B) leads to gains in structure-related metrics, including $S_{\alpha}$ and $E_{\phi}$, indicating that AoLP provides useful directional and structural information for camouflage perception. Introducing DoLP alone (C) further improves the results, suggesting that DoLP also provides effective complementary cues for distinguishing camouflaged targets. These observations indicate that both AoLP and DoLP are beneficial, while capturing different aspects of polarization information. However, each cue alone can only provide partial polarization evidence. This motivates our design to jointly model them for constructing more reliable polarization guidance, and supports our design philosophy of treating polarization as structural guidance rather than direct feature fusion.
@@
Effectiveness of the Iterative Feedback Decoder. Finally, removing the Iterative Feedback Decoder (IFD) (I) results in the most noticeable performance drop, especially in $F_{\beta}^{w}$ and $S_{\alpha}$, indicating that coarse-to-fine feedback refinement is crucial for recovering ambiguous structures and improving boundary completeness. As shown in Fig. 6, the qualitative results further support our findings.
Overall Discussion. Overall, all ablation variants perform worse than the full model, which confirms that the superiority of CPGNet comes from the coordinated effect of PIM, PGF, and IFD, where PGF further benefits from the joint contribution of PGE and EFM. This also suggests that effective camouflage perception requires not only reliable polarization guidance, but also its progressive propagation and refinement throughout the network.
@@ 5. Limitation
The proposed CPGNet still has several limitations. Its performance depends on the quality of polarization measurements, and unstable DoLP/AoLP responses in weak-polarization regions may affect the reliability of the derived guidance cues. Fig. 7 shows that RGB lacks discriminative features, while both polarization and depth modalities offer limited complementary cues, making it challenging for many methods to detect camouflaged targets.